S(s)+KNO3(s)+7C(s) ---> K2S(s)+K2CO3(s)+2N2(g)+3CO(g)+3CO2(g) Δ= -1688,2 kJ
From the cartridge existing gunpowder during combustion there have been released 21,1 kJ heat. Measure how much carbon was in cartridge.
1
Expert's answer
2019-11-26T06:40:19-0500

Solution.

Let's make a proportion:

"\frac{n(C)}{7} = \frac{-21.1}{-1688.2}"

n(C) = 0.09 mol

"n(C) = \frac{m(C)}{M(C)}"

m(C) = 0.09 * 12 = 1.08 g

Answer:

n(C) = 0.09 mol

m(C) = 1.08 g